OpenStudy (destroyer5):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −2 x = 0 x = 2 There is no x value, as there is no solution. - This one x can be any value, as there is an infinite number of solutions.

Vocaloid (vocaloid):

hint: multiply the first equation by 2

OpenStudy (destroyer5):

ok thanks

Vocaloid (vocaloid):

let me know if you get stuck

OpenStudy (destroyer5):

i got 2x+y=2, so x would be 2? right

Vocaloid (vocaloid):

nope

OpenStudy (destroyer5):

i substituted what i needed to?

Vocaloid (vocaloid):

what is 2x * 2?

OpenStudy (destroyer5):

4x

Vocaloid (vocaloid):

good, what is y * 2?

OpenStudy (destroyer5):

isnt it 2y

Vocaloid (vocaloid):

good so that gives us: 4x + 2y = 2 4x + 2y = -2

Vocaloid (vocaloid):

now subtract the two equations let me know what you get

OpenStudy (destroyer5):

x= 1/4 and y=-3/2

Vocaloid (vocaloid):

??????

OpenStudy (destroyer5):

no?

Vocaloid (vocaloid):

what is 4x - 4x?

OpenStudy (destroyer5):

0 of course

Vocaloid (vocaloid):

good now what is 2y - 2y?

OpenStudy (destroyer5):

0

Vocaloid (vocaloid):

good what about 2 - (-2)?

OpenStudy (destroyer5):

4?

Vocaloid (vocaloid):

good! so that gives us: 0 = -4 which has no solution

OpenStudy (destroyer5):

ohh ok thank you so much i did consider that but wasnt sure!

OpenStudy (destroyer5):

so its c!

OpenStudy (q-tip):

Vocaloid has just given you the elimination method. Here is another method but it's by substitution (which is also what the question wants) 2x + y = 1 4x + 2y = −2 Rearrange the first equation so that y is on one side: y=1-2x That is the value for y, so you can plug that in for your second equation: 4x + 2(1-2x) = -2 Expand: 4x+2-4x=-2 Simplify: 2=-2 Obviously this is incorrect so the answer is C.